# Characterizing Hitchin representations over real closed fields

### v2

@@ Abstract.
The goal of this article is to show that Hitchin representations over real closed field extensions $\mathbb{F}$ of $\mathbb{R}$ are exactly those representations from the fundamental group of a closed surface into $\mbox{\rm PSL}(n,\mathbb{F})$ that are conjugate to $\mathbb{F}$-positive representations, i.e. representations that admit an equivariant limit map from the set of fixed points in the boundary of the universal cover of the surface into the set of full flags in $\mathbb{F}^{n}$ satisfying specific positivity properties. They form semi-algebraically connected components of the space of all representations that consist entirely of injective representations with discrete image.
@@ 1. Introduction
Let $S$ be a closed, connected and oriented surface of genus at least two. Higher Teichmüller theory concerns itself with the study of connected components of the character variety $\chi(S,G)$, for $G$ a semisimple Lie group of higher rank, consisting entirely of injective representations with discrete image, called higher Teichmüller spaces [Wie18]. They are conjectured to parametrize geometric structures on $S$, see e.g. [CG93, GW08]. To understand these and degenerations thereof it is often useful to compactify them in a geometrically meaningful way. In this article, we concentrate on the real spectrum compactification of $\chi(S,G)$, whose study was initiated by Burger–Iozzi–Parreau–Pozzetti in their seminal works [BIPP21, BIPP23], inspired by the work of Brumfiel in the case of Teichmüller space [Bru88]. Its boundary points are identified with equivalence classes of reductive representations from $\pi_{1}(S)$ into $G_{\mathbb{F}}$, where $\mathbb{F}$ is a real closed field containing $\mathbb{R}$ and $G_{\mathbb{F}}$ is the $\mathbb{F}$-extension of $G$ (2.6) [BIPP23, Theorem 1.1]. For the purpose of this article this will serve as a definition. Thus boundary points can be studied in terms of representations. For a higher Teichmüller space $\mathcal{C}$ of $\chi(S,G)$ a natural question in this context is the following: Given a reductive representation from $\pi_{1}(S)$ to $G_{\mathbb{F}}$, can we determine “geometrically” whether or not it is in the boundary of $\mathcal{C}$?
We answer this question in the case when $\mathcal{C}$ is the Hitchin component—a higher Teichmüller space of the $\mbox{\rm PSL}(n,\mathbb{R})$-character variety. In analogy to the real case, see [FG06, Lab06, Gui08], we give a characterization of $\mathbb{F}$-Hitchin representations in terms of positivity of an equivariant limit map from a subset of the boundary of the universal cover of $S$ to the set of full flags in $\mathbb{F}^{n}$.
@@
For maximal components in $\chi(S,\textrm{Sp}(2n,\mathbb{R}))$ the study of the above question was announced in [BIPP21]. Similarly, one can ask if the closure of a connected component consisting of $\Theta$-positive representations in the real spectrum compactification of $\chi(S,G)$, for $G$ a Lie group admitting a $\Theta$-positive structure, can be characterized in a similar way [GW18, GLW21, GW22, BIPP23].
Let us explain the objects and how the main result provides an answer to the above question more precisely. In his seminal paper [Hit92], Hitchin used Higgs bundles to show that $\chi(S,\mbox{\rm PSL}(n,\mathbb{R}))$ for $n\geq 3$ contains three connected components, if $n$ is odd, and six connected components, if $n$ is even. In the odd case, one of the three components, and in the even case, two of the six components, are homeomorphic to $\mathbb{R}^{(2g-2)(n^{2}-1)}$. They can be characterized as follows. Denote by $\iota_{n}$ the irreducible $n$-dimensional representation of $\mbox{\rm PSL}(2,\mathbb{R})$ into $\mbox{\rm PSL}(n,\mathbb{R})$.
@@
From the construction of $\textnormal{RSp}(\chi(S,\mbox{\rm PSL}(n,\mathbb{R})))$ it follows that $\mathbb{F}$-Hitchin representations are exactly those representations representing points in the boundary of $\textnormal{Hit}(S,n)$ in the real spectrum compactification of the character variety, see e.g. [BCR98, Chapter 7]. In fact, the following corollary of the above considerations and 1.1 completes the proof of a result announced in [BIPP21, Theorem 46] and answers the question of geometric characterization for $\mathcal{C}=\textnormal{Hit}(S,n)$. Denote by $\textnormal{Hit}(\bar{S},n)$ the connected component of $\chi(S,\mbox{\rm PSL}(n,\mathbb{R}))$ containing $\iota^{\prime}_{n}\circ j$, where $\iota^{\prime}_{n}$ is any conjugate of $\iota_{n}$ under an element of ${\rm PGL}(n,\mathbb{R})\setminus\mbox{\rm PSL}(n,\mathbb{R})$. If $n$ is odd, $\textnormal{Hit}(\bar{S},n)$ is the same as $\textnormal{Hit}(S,n)$.
@@
We finish by proving general properties for $\mathbb{F}$-Hitchin representations.
@@
Let $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ be $\mathbb{F}$-Hitchin. Then $\rho$ is injective, discrete, positively hyperbolic (6.1) and dynamics-preserving (6.3).
@@
Method
The key tool to establish the forward direction of 1.1 is the Tarski–Seidenberg transfer principle (2.7), a real closed analogue of the Lefschetz principle for algebraically closed fields. It enables to deduce results for real closed extensions of $\mathbb{R}$, that are stated in first-order logic in the language of ordered fields, from results that hold true over $\mathbb{R}$. The transfer principle proves positive hyperbolicity for $\mathbb{F}$-Hitchin representations. Equipped with this we construct a limit map for an $\mathbb{F}$-Hitchin representation $\rho$. Namely to the fixed point of a hyperbolic element $\gamma$ we associate the stable flag of the positively hyperbolic element $\rho(\gamma)$. To prove its positivity properties, we use again the Tarski–Seidenberg transfer principle and the positivity properties of real Hitchin representations (7.1).
@@
The backward direction of 1.1 is more involved. Here we cannot apply the transfer principle, as the notion of $\mathbb{F}$-positive representation involves infinitely many conditions. We will show that, actually, a finite set of data points suffices to determine an $\mathbb{F}$-positive representation up to conjugation. The equivariant limit map that comes together with an $\mathbb{F}$-positive representation allows us to associate to the representation the variant of the Bonahon–Dreyer coordinates over $\mathbb{F}$. We then verify that these coordinates satisfy the same polynomial equalities and inequalities as the ones for $\mathbb{F}$-Hitchin representations, and that two $\mathbb{F}$-positive representations have the same coordinates if and only if they are conjugate. This proves 1.6, and then also 1.1. The main difficulty compared to the case $\mathbb{F}=\mathbb{R}$ lies in the proof of the closed leaf inequality (iv) in Section 7.2. For this we prove that $\mathbb{F}$-positive representations are positively hyperbolic and dynamics-preserving (6.4) without any analytic tools.
@@
The article is organized as follows. Section 2 states the important notions and results from real algebraic geometry and sets up notation. We introduce Hitchin representations over real closed fields in Section 3. In Section 4 we give the necessary preliminaries on flags, including the definition of the triple and double ratios. This leads to the notions of positive tuples of full flags in Section 5, and we hint at the connection between positivity of flags and positivity properties of matrices. Even though most results from Section 2-5 are known to the experts for $\mathbb{R}$, we carefully study all objects in question focussing in particular on their semi-algebraic structures, with the goal of generalizing to real closed fields different from $\mathbb{R}$.
@@
Section 9 collects properties of $\mathbb{F}$-Hitchin representations and finishes the proof of 1.7.
Acknowledgements
I thank Marc Burger for introducing me to this topic and his helpful remarks. A special thanks goes to Giuseppe Martone for pointing me to the correct way of proving 6.4. I am particularly grateful to Jacques Audibert, Karin Baur, Peter Feller, Gerhard Hiß, Alessandra Iozzi, Fanny Kassel, Mareike Pfeil, Valdo Tatitscheff and Tengren Zhang for constructive discussions and detailed feedback on this manuscript.

### v3

# Real spectrum compactification of Hitchin components, Weyl chamber valued lengths, and dual spaces

@@ Abstract.
The main result of this article is that Hitchin representations over real closed field extensions $\mathbb{F}$ of $\mathbb{R}$ correspond precisely to those representations of the fundamental group of a closed surface into $\mbox{\rm PSL}(n,\mathbb{F})$ that are conjugate to $\mathbb{F}$-positive representations, i.e. representations that admit an equivariant limit map from the set of fixed points in the boundary of the universal cover of the surface into the set of full flags in $\mathbb{F}^{n}$ satisfying specific positivity properties. As the theorem treats general real closed fields, and not only the reals, the tools of analysis are not available. Instead, our proof is based on the Tarski–Seidenberg transfer principle and a multiplicative version of the Bonahon–Dreyer coordinates.
We use this result to prove that $\mathbb{F}$-positive representations form semi-algebraically connected components of the space of all representations, that consist entirely of injective and discrete representations, which are positively hyperbolic and weakly dynamics-preserving over $\mathbb{F}$. Furthermore, we show how to associate intersection geodesic currents to $\mathbb{F}$-positive representations, and conclude with applications to the Weyl chamber length compactification and to dual spaces of geodesic currents.
Key words and phrases:
higher Teichmüller theory, real algebraic geometry, compactifications
1991 Mathematics Subject Classification:
22E40, 30F60, 14P10
@@ 1. Introduction
Let $S$ be a closed, connected and oriented surface of genus at least two. Higher Teichmüller theory concerns itself with the study of connected components of the character variety $\chi(S,G)$, for $G$ a semisimple Lie group of higher rank, consisting entirely of injective representations with discrete image, called higher Teichmüller spaces [Wie18]. They are conjectured to parametrize geometric structures on $S$; see e.g. [CG93, GW08]. To understand these and degenerations thereof it is often useful to compactify them in a geometrically meaningful way; see for example [MS84, Bon88, Lof04, Ale07, Par12, OT21, LTW22, SS22, Rei23].
In this article, we concentrate on the real spectrum compactification of $\chi(S,G)$, whose study was initiated by Burger–Iozzi–Parreau–Pozzetti in their seminal works [BIPP21b, BIPP23], inspired by the work of Brumfiel in the case of Teichmüller space [Bru88]. Its boundary points are identified with equivalence classes of reductive representations from $\pi_{1}(S)$ into $G_{\mathbb{F}}$, where $\mathbb{F}$ is a real closed field containing $\mathbb{R}$ and $G_{\mathbb{F}}$ is the $\mathbb{F}$-extension of $G$ (2.6) [BIPP23, Theorem 1.1]. Thus boundary points can be studied in terms of representations. For a higher Teichmüller space $\mathcal{C}$ of $\chi(S,G)$ a natural question in this context is the following: Given a reductive representation from $\pi_{1}(S)$ to $G_{\mathbb{F}}$, can we determine “geometrically” whether or not it is in the boundary of $\mathcal{C}$?
We answer this question in the case when $\mathcal{C}$ is the Hitchin component—a higher Teichmüller space of the $\mbox{\rm PSL}(n,\mathbb{R})$-character variety. In analogy to the real case, see [FG06], we give a characterization of $\mathbb{F}$-Hitchin representations in terms of positivity of an equivariant limit map from a subset of the boundary of the universal cover of $S$ to the set of full flags in $\mathbb{F}^{n}$.
@@
This characterization gives a semi-algebraic description of the set of $\mathbb{F}$-positive representations and thus allows a direct study of these representations via Hitchin representations. For example, our description allows to prove that for every $\mathbb{F}$-positive representation $\rho$ there exists a continuous path $\rho_{t}$ of real Hitchin representations, such that the renormalized length functions of $\rho_{t}$ converge to the length function of $\rho$ (1.11). As the theorem treats general $\mathbb{F}$ instead of just the reals, the tools of analysis—crucially used for the analog result over $\mathbb{R}$—are not available. Instead, our proof is based on carefully describing and characterizing all involved notions semi-algebraically and using the Tarski–Seidenberg transfer principle. Furthermore, we provide a multiplicative variant of the Bonahon–Dreyer-coordinates for the set of $\mathbb{F}$-Hitchin representations. For maximal components in $\chi(S,\textrm{Sp}(2n,\mathbb{R}))$ the analogue of the above result was announced in [BIPP21b]. More generally, one can ask if the closure of the set of $\Theta$-positive representations in the real spectrum compactification of $\chi(S,G)$, for $G$ a Lie group admitting a $\Theta$-positive structure, can be characterized in a similar way [GW18, GLW21, GW25, BIPP23].
Let us now explain the objects and results of this article in more detail. In his seminal paper [Hit92], Hitchin used Higgs bundles to show that $\chi(S,\mbox{\rm PSL}(n,\mathbb{R}))$ for $n\geq 3$ contains three connected components, if $n$ is odd, and six connected components, if $n$ is even. In the odd case, one of the three components, and in the even case, two of the six components, are homeomorphic to $\mathbb{R}^{(2g-2)(n^{2}-1)}$. They can be characterized as follows. Denote by $\iota_{n}$ the irreducible $n$-dimensional representation of $\mbox{\rm PSL}(2,\mathbb{R})$ into $\mbox{\rm PSL}(n,\mathbb{R})$.
@@
From the construction of $\textnormal{RSp}(\chi(S,\mbox{\rm PSL}(n,\mathbb{R})))$ it follows that $\mathbb{F}$-Hitchin representations are exactly those representations representing points in the boundary of $\textnormal{Hit}(S,n)$ in the real spectrum compactification of the character variety, see Section 3.4. In fact, the following corollary of the above considerations and 1.1 completes the proof of a result announced in [BIPP21b, Theorem 46] and answers the question of geometric characterization for $\mathcal{C}=\textnormal{Hit}(S,n)$.
Denote by $\textnormal{Hit}(\bar{S},n)$ the connected component of $\chi(S,\mbox{\rm PSL}(n,\mathbb{R}))$ containing $\iota^{\prime}_{n}\circ j$, where $\iota^{\prime}_{n}$ is any conjugate of $\iota_{n}$ under an element of ${\rm PGL}(n,\mathbb{R})\setminus\mbox{\rm PSL}(n,\mathbb{R})$. If $n$ is odd, $\textnormal{Hit}(\bar{S},n)$ is the same as $\textnormal{Hit}(S,n)$.
@@
We summarize general properties for $\mathbb{F}$-Hitchin representations.
@@
Let $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ be $\mathbb{F}$-Hitchin. Then $\rho$ is injective, discrete [^1] Careful, unless $\mathbb{F}=\mathbb{R}$ the topology is not locally compact., positively hyperbolic (6.1) and weakly dynamics-preserving (6.3).
@@
Next we discuss how to associate a geodesic current to an $\mathbb{F}$-positive representation. Geodesic currents were introduced by Bonahon in [Bon86], providing a unifying framework for Thurston’s compactification. They are now an active field of research by themselves and have applications in various fields in the study of surfaces; see e.g. [ES22]. To simplify we choose a hyperbolic structure on $S$, even though the following concept can be defined in a purely topological way, see e.g. [Bon88, Fact 1].
Definition 1.8.
A geodesic current on $S$ is a locally finite, $\pi_{1}(S)$-invariant, regular Borel measure on the space of (unoriented and unparametrized) geodesics in the universal cover $\tilde{S}$ of $S$.
Denote by $\mathscr{C}(S)$ the space of geodesic currents on $S$ endowed with the weak-topology, and by $\mathbb{P}\mathscr{C}(S):=\mathscr{C}(S)/\mathbb{R}_{>0}$ the space of projectivized geodesic currents, the latter being compact [Bon88, Proposition 4, Corollary 5]. Many seemingly different objects are in fact geodesic currents, e.g. homotopy classes of closed curves on $S$ or isotopy classes of marked hyperbolic structures on $S$ [Bon88, Lemma 9]. Martone-Zhang proved in [MZ19, Theorem 3.4], that we can associate to a Hitchin representation a geodesic current such that $k$-length functions of the representation can be computed as intersections (10.1) with this current. More precisely, let $\mathbb{F}\supseteq\mathbb{R}$ be a real closed field that admits an order-compatible valuation $\upsilon$, see 10.4 (for $\mathbb{F}=\mathbb{R}$ we can take $\upsilon=-\log$). For $g\in\mbox{\rm PSL}(n,\mathbb{F})$ with distinct eigenvalues of the same sign, let $\lambda_{1}(g)>\ldots>\lambda_{n}(g)>0$ be the eigenvalues of a lift of $g$ to $\mbox{\rm SL}(n,\mathbb{F})$. For every $k=1,\ldots,n-1$ we define the $k$-length of $g$ as
$L_{k}(\rho(\gamma))\coloneqq-\sum_{j=1}^{k}\upsilon(\lambda_{j}(g))+\sum_{j=n-k+1}^{n}\upsilon(\lambda_{j}(g)).$ (1)
These length functions can be interpreted as the translation length of an element $g\in\mbox{\rm PSL}(n,\mathbb{F})$ acting on the metric space $\mathscr{B}_{\textnormal{PSL}(n,\mathbb{F})}$ as defined in [BIPP21b, Section 3.4 and 4]—the non-Archimedean analogue of the symmetric space of $\mbox{\rm PSL}(n,\mathbb{R})$ that generalizes real trees to higher rank.
We establish the following result announced by Burger–Iozzi–Parreau–Pozzetti [BIPP21b, Theorem 47]. An equivalent version was proven for maximal representations by Burger–Iozzi–Parreau–Pozzetti [BIPP21a, Theorem 1.2], and recently extended to positively ratioed representations [BIPP23, Theorem 9.6]. Recall that to every $\rho\in\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(n,\mathbb{R}))$ are associated $n-1$ geodesic currents [MZ19]. This result can be extended in the following way.
Theorem 1.9.
Let $\mathbb{F}\supseteq\mathbb{R}$ be a real closed field with an order-compatible valuation $\upsilon$ (assumed to be $-\log$ if $\mathbb{F}=\mathbb{R}$) and let $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ be $\mathbb{F}$-positive. Then for every $k=1,\ldots,n-1$, there exists a geodesic current $\mu^{k}_{\rho}$ such that for any $e\neq\gamma\in\pi_{1}(S)$ we have
$i(\mu^{k}_{\rho},\gamma)=L_{k}(\rho(\gamma)).$
The current $\mu_{\rho}^{k}$ is non-zero if and only if there exists $\gamma\in\pi_{1}(S)$ with $\upsilon(|\mbox{\rm tr}(\rho(\gamma))|)<0$.
Considering the subset $\textnormal{RSp}^{\textnormal{cl}}(\textnormal{Hit}(S,n))$ of closed points in $\textnormal{RSp}(\textnormal{Hit}(S,n))$, this implies the following.
Corollary 1.10.
For all $k=1,\ldots,n-1$ the map
$\textnormal{RSp}^{\textnormal{cl}}(\textnormal{Hit}(S,n))$ $\to\mathbb{P}\mathscr{C}(S),$
$[(\rho,\mathbb{F})]$ $\mapsto\left[\mu^{k}_{\rho}\right]$
is well-defined and continuous.
Applications
We end the introduction with two applications of our results: one concerning Weyl chamber length functions [Par12], and the other dual spaces of geodesic currents [DRMG22]. The objects will be introduced in detail in Section 11.
Let $\mathbb{F}$ be a real closed field with an order-compatible valuation $\upsilon\colon\mathbb{F}\to\mathbb{R}\cup\{\infty\}$ (if $\mathbb{F}=\mathbb{R}$, we take $\upsilon=-\log$). Given $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ an $\mathbb{F}$-positive representation, then $\rho(\gamma)$ is positively hyperbolic with eigenvalues $\lambda_{1}(\rho(\gamma))>\ldots>\lambda_{n}(\rho(\gamma))>0$. This is classical for $\mathbb{F}=\mathbb{R}$ and follows from 6.4 for general $\mathbb{F}$. Define the Weyl chamber length function
$L_{\rho}\colon\pi_{1}(S)$ $\to\mathfrak{a}^{+},$
$\gamma$ $\mapsto(-\upsilon(\lambda_{1}(\rho(\gamma))),\ldots,-\upsilon(\lambda_{n}(\rho(\gamma)))),$
where $\mathfrak{a}$ denotes the Cartan subalgebra of the Lie algebra of $\mbox{\rm SL}(n,\mathbb{R})$ consisting of diagonal matrices of trace zero, and $\mathfrak{a}^{+}$ the Weyl chamber consisting of those elements $\textnormal{diag}(a_{1},\ldots,a_{n})\in\mathfrak{a}$ with $a_{1}>\ldots>a_{n}$.
Corollary 1.11.
Let $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ be an $\mathbb{F}$-positive representation, where $\mathbb{F}$ is a non-Archimedean field with an order-compatible valuation. Assume $L_{\rho}$ does not vanish identically. Then there is a continuous piecewise semi-algebraic path
$\mathbb{R}_{\geq 0}\to\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(n,\mathbb{R})),\quad t\mapsto\rho_{t}$
and a constant $c>0$ such that
$\lim_{t\to\infty}\frac{L_{\rho_{t}}(\gamma)}{t}=cL_{\rho}(\gamma).$
In particular, the renormalized $k$-length functions (as defined in (1)) converge (up to scaling) to the length function of $\rho$.
This has the following application to dual spaces of geodesic currents. Recall that to any current $\mu$ one associates the dual space as defined in [DRMG22]. In the case when $\mu$ is the geodesic current associated to some $\rho\in\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(3,\mathbb{R}))$, taking the dual space gives a distance on $\mathbb{H}^{2}$ that is isometric $\Omega_{\rho}$ endowed with its Hilbert metric, where $\Omega_{\rho}$ is the properly convex subset of $\mathbb{R}\mathbb{P}^{2}$ associated to $\rho$ [CG93]. Using the above corollary we obtain the following.
Corollary 1.12.
In the setting of 1.11 for $n=3$, the sequence of convex projective structures with Hilbert metrics $(\Omega_{\rho_{t}},\tfrac{1}{t}d_{\Omega_{\rho_{t}}})$ converges in the equivariant Gromov–Hausdorff topology to the dual space of the current $\mu_{\rho}$ associated to $\rho$ (up to scaling).
Methods
@@
The key tool to establish the forward direction of 1.1 is the Tarski–Seidenberg transfer principle (2.7), a real closed analogue of the Lefschetz principle for algebraically closed fields. It enables to deduce results for real closed extensions of $\mathbb{R}$, that are stated in first-order logic in the language of ordered fields, from results that hold true over $\mathbb{R}$. This allows to prove positive hyperbolicity for $\mathbb{F}$-Hitchin representations. Equipped with this we construct a limit map for an $\mathbb{F}$-Hitchin representation $\rho$. Namely to the fixed point of a hyperbolic element $\gamma$ we associate the stable flag of the positively hyperbolic element $\rho(\gamma)$. To prove its positivity properties, we use again the Tarski–Seidenberg transfer principle and the positivity properties of real Hitchin representations (7.1). The backward direction of 1.1 is more involved. Here we cannot apply the transfer principle, as the notion of $\mathbb{F}$-positive representation involves infinitely many conditions. We will show that, actually, a finite set of data points suffices to determine an $\mathbb{F}$-positive representation up to conjugation. The equivariant limit map that comes together with an $\mathbb{F}$-positive representation allows us to associate to the representation the variant of the Bonahon–Dreyer coordinates over $\mathbb{F}$. We then verify that these coordinates satisfy the same polynomial equalities and inequalities as the ones for $\mathbb{F}$-Hitchin representations, and that two $\mathbb{F}$-positive representations have the same coordinates if and only if they are conjugate. This proves 1.6, and then also 1.1. The main difficulty compared to the case $\mathbb{F}=\mathbb{R}$ lies in the proof of the closed leaf inequality (iv) in Section 7.2. For this we prove that $\mathbb{F}$-positive representations are positively hyperbolic and weakly dynamics-preserving (6.4) without any analytic tools, relying solely on their positivity properties.
To prove 1.9 we show that $\mathbb{F}$-positive representations are positively ratioed using the Tarski–Seidenberg transfer principle, to then apply a result of Burger–Iozzi–Parreau–Pozzetti about geodesic currents associated to positive cross-ratios (10.3).
@@
The article is organized as follows. Section 2 states the important notions and results from real algebraic geometry and sets up notation. We introduce Hitchin representations over real closed fields in Section 3. In Section 4 we give the necessary preliminaries on flags, including the definition of the triple and double ratios. This leads to the notion of positive tuples of full flags in Section 5, and we hint at the connection between positivity of flags and positivity properties of matrices. Even though most results from Section 2-5 are known to the experts for $\mathbb{R}$, we carefully study all objects in question focusing in particular on their semi-algebraic structures, with the goal of generalizing to real closed fields different from $\mathbb{R}$.
@@
Section 9 collects properties of $\mathbb{F}$-Hitchin representations and finishes the proof of 1.7. In Section 10, we give some background on geodesic currents, positive cross-ratios, and valuations. We then prove 1.9 and 1.10. The proof of the two applications in Section 11 finishes the paper.
Acknowledgments
I thank Marc Burger for introducing me to this topic, his helpful remarks and explaining the interesting applications of the main result to me. A special thanks goes to Giuseppe Martone for pointing me to the correct way of proving 6.4. I am particularly grateful to Jacques Audibert, Karin Baur, Peter Feller, Gerhard Hiß, Alessandra Iozzi, Fanny Kassel, Mareike Pfeil, Valdo Tatitscheff and Tengren Zhang for constructive discussions or detailed feedback on this manuscript.